Tenecteplase is a modified enzyme that converts plasminogen to plasmin, which aids in the breakdown of blood clots. The recommended dose of tenecteplase for a 165-pound patient is 40 mg, which is administered by IV bolud iver a 5-second interval. The concentrwation of the tenexteplase solution is 5 mg/ml. How naby milliters are present in a dingle dose of the solution?
Answer #1

Ordered dose for 165 pound patient=40 mg

1ml solution = 5 mg

So, 40 mg= X ml

Dose in ml=( 1/5= X/40)

X= 40/5= 8 ml